OpenStudy (anonymous):

What is the slope of the line through the points (2, 6) and (3, 5)? m = _?_ (Write only the value of the slope in the answer blank, do not use any variables.) Numerical Answers Expected!

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

5-6/3-2= -1

OpenStudy (jwheele1):

y1 = 6 y2 = 5 x1 = 2 x2 = 3

OpenStudy (jwheele1):

5-6 = -1 3-2 = 1

OpenStudy (jwheele1):

\[\frac{ -1 }{ 1 } = -1\]

OpenStudy (jwheele1):

answer = -1

OpenStudy (anonymous):

it's -1 for the answer u guys r talkin bout

OpenStudy (anonymous):

i think @rebecca1233 has got her answer or is she playin lol

OpenStudy (jwheele1):

Well, I was showing my work so that rebecca could learn how to do it herself better.

OpenStudy (anonymous):

thanks (: